OpenStudy (anonymous):

Evaluate: if g(x)=3x^4/3, find the value of g(-27)

OpenStudy (mertsj):

What is the third root of -27?

OpenStudy (mertsj):

In other words what makes this sentence true: ( ? )^3=-27

OpenStudy (anonymous):

3 @Mertsj

OpenStudy (mertsj):

3x3x3=27 not -27

OpenStudy (anonymous):

-3?

OpenStudy (mertsj):

And what is (-3)^4

OpenStudy (anonymous):

-81

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

And multiply that by 3

OpenStudy (anonymous):

-243

OpenStudy (mertsj):

\[3(-27)^{\frac{4}{3}}=3((\sqrt[3]{-27}))^4=3(-3)^4=3(81)=243\]